Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day06


Crew #108 Log Book for December 08, 2011

Photos of the day





Commander's Check-In Report

Date and time: 

8 December 2011


Written by : 

Charlotte Poupon, Commander


Crew Physical Status:


The sick crew member is out of sim due to recurrent kidney infection and is about to fly back with a cinema crew.

Time Departed/Returned from EVA:
EVA # 11 / 11:00 to 12:00
EVA # 12 / 15:00 to 16:50

Brief Narrative of Field Mission Results:

Not applicable


EVA Data/Interpretations:

The first EVA went to Phobos Peak.
Geologist found a lots of concretions.


Engineering/Hab Maintenance:

The power failure declared sooner was due to John changing stuff on the generator. So doesn't have to be considered as a failure… even if the question about HOW to warn mission control in case of real failure remains. No electricity = no wifi = no internet.


Report Transmission Schedule:

All reports by 22:00.


Plans for Tomorrow:

Light installation is on. Hourly measurement should start by 07:00


Inventory:

Not applicable


Support Requested:

Not applicable


Miscellaneous: 

N/A

Commander's Report

Date and time: 

8 December 2011


Written by : 

Charlotte Poupon, Commander


The day was marked by one crew member's anticipated departure for health reasons, as we mentioned in our latest reports.

The film crew did not need to film more so after a good-bye visit they left the 5 of us to complete the rotation.

Station leader will assume shifts on duty regarding tidying and cooking, in order to not overload other crew members that are quite busy, especially the engineers.

It was the second non-cooking day and generally crew members have no
specific complaints about that. Tomorrow opens a shift of 2 cooking days.


Engineering Report

Date and time: 

8 December 2011


Written by: 

Ashley Dale, Crew 108, Engineer


Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes
Notes/Comments (include how many generator runs, approximately when,
and any times you needed to turn the generator on/off): Don came and
changed the oil for the generator


Oil:
Kitty Oil Quantity (# of Quarts in storage): Unknown
Honey Oil Quantity (# of Quarts in storage): Unknown
ATV Oil Quantity (# of Quarts in storage): 3
Car Oil Quantity (# of Quarts in storage): 3

Water Status:
Outside Potable Water Tank Level (inches from bottom): 0.5
Trailer Potable Water Tank Level (inches from bottom): 18
Hab Potable Water Tank Level (inches from bottom): Full

Observatory:
Telescope Used? (Yes/No): No
Observatory Notes/Comments: Mission Support are trying to contact the
the relevant person to unlock it for use

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no):No
Tire Status: Good
Notes/Comments on Hab Car: N/A

ATV 1:
Used: Yes
Oil Checked: Yes
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: Oil was taken out and replaced

ATV 2:
Used: Yes
Oil Checked: Yes
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: Oil was taken out and replaced

ATV 3:
Used: Yes
Oil Checked: Yes
Fuel Consumed: 0.2 gallons
Tire Status: Good
Comments: Oil was taken out and replaced

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 66.7
Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): Unknown

Computers/Networking Infrastructure:
Notes/comments on internet/computers?: the VisionGS, the http port is
set to 80 - but nothing is uploading

General Engineering Notes/Comments: 
EVA suits 5 and 6 required repairing, tempermental securing of battery pack on EVA suit 2 has been improved, rattling fan has been cleaned and quietened.

Journalist Report

Date and time: 

8 December 2011, 21:49


Written by: 

Ashley Dale, Crew 108, Engineer

Caught another really great sunrise. The film crew arrived this morning with news that one crew member,  had to leave because of a recurring kidney related issue – so we are down to a crew of five. Filming was completed so hopefully the art film produced will bring something worthwhile to the Mars Society and its cause. The engineers carried out general maintenance on the EVA suits, and made a few minor repairs around the Hab. There were a couple of exploratory EVAs, a few fossils were found, some petrified wood, and an area of scientific interest was located at Phoebes Peak for our astrobiologist - she will return in an EVA tomorrow. The Mars Rover our commander has completed putting together, BuddyBot, will have a trial run around the Hab area tomorrow before we decide to send it down any of the caves/canyons located.

EVA 10 Report


Written by:

Ashley Dale, Crew 108, Engineer & Hab Support for EVA 10 and Mike Lotto, Crew 108, Engineer


Time: 

09:14 – 10:59

- Depressurization: 09h14 – 09h19

- Reached based of Phoebes Peak: 09h42

- Reach top of Phoebes Peak: 09h57

- Descent: 10h01 – 10h07

- ATV to Caves: 10h07 – 10h018

- Reached peak of caves: 10h31

- Emergency Repair (Lotto) on EVA backpack 2 (Lingappa): 10h32 – 10h38

- Descent: 10h43 – 10h49

- Hab Support (Dale) recommends descent route: 10:46

- ATV to Hab: 10h49 – 10h53

- Repressurization: 10h54 – 10h59

Crew Members: 

Michael LeClair (Geophysicist), Usha Lingappa (Astrobiologist), Mike Lotto (Engineer)


Site Location:

Phoebes Peak, Olympus Mons


Transit Mode: 

 ATV


Objectives:

Explore regions for Mars Rover use, explore locations of geophysical interest, explore Phoebes Peak to the east of the Hab


Results: Rock varnish found on Phoebes Peak (N 51:93:26, W 42:50:371) – will return with equipment on later EVA for Lingappa. Nothing of interest in rock caverns on plateau of Olympus Mons (N 51:84:88, W 42:51:667).

Lessons Learned: 

Plateau of Olympus Mons has nice views but nothing else. Phoebes Peak is a good site for astrobiology. Battery on EVA Suit 2 needs to be secured.

EVA 12 Report

Date and time: 

8 December 2011

Written by: 

Michael LeClair, Crew 108, XO

Time: 

15:02 - 16:48

Crew Members: 

Michael LeClair (Geologist), Ashley Dale (Engineer), Lotto (Engineer)

Site Location: 

Ridge line

Transit Mode: 

ATV

Objectives: 

Collect concretions samples

Results: 

Objectives attained; 4 samples collected

Lessons Learned: 

Ample samples along ridgeline